# Deceptive Sequential Decision-Making via Regularized Policy Optimization

### v2

@@ Abstract
Autonomous systems are increasingly expected to operate in the presence of adversaries, though adversaries may infer sensitive information simply by observing a system. Therefore, present a deceptive sequential decision-making framework that not only conceals sensitive information, but actively misleads adversaries about it. We model autonomous systems as Markov decision processes, with adversaries using inverse reinforcement learning to recover reward functions. To counter them, we present three regularization strategies for policy synthesis problems that actively deceive an adversary about a system’s reward. “Diversionary deception” leads an adversary to draw any false conclusion about the system’s reward function. “Targeted deception” leads an adversary to draw a specific false conclusion about the system’s reward function. “Equivocal deception” leads an adversary to infer that the real reward and a false reward both explain the system’s behavior. We show how each form of deception can be implemented in policy optimization problems and analytically bound the loss in total accumulated reward induced by deception. Next, we evaluate these developments in a multi-agent setting. We show that diversionary, targeted, and equivocal deception all steer the adversary to false beliefs while still attaining a total accumulated reward that is at least $97\%$ of its optimal, non-deceptive value.
Index Terms:
Deception, Markov Decision Processes, Inverse Reinforcement Learning
@@ I Introduction
Existing work in [18] has presented related definitions of deception and empirical results for them, but to the best of our knowledge the current paper is the first to analytically characterize these types of deception and the tradeoffs they create.
@@ II-A Preliminaries on Markov Decision Processes
$\begin{aligned} \textnormal{ subject to }\,\,&v_{\pi}(s)-\gamma\sum_{j\in\mathcal{S}}\mathcal{T}(s,a,j)v_{\pi}(j)\geq r(s,a)\quad\textnormal{ for all }s\in\mathcal{S},a\in\mathcal{A}.\end{aligned}$
@@
$\begin{aligned} \underset{x\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}}{\operatorname{maximize}}\quad&\sum_{s\in S}\sum_{a\in A}r(s,a)x(s,a)\end{aligned}$ (3)
@@ II-B1 Apprenticeship Learning
  $\bar{\mu}^{(i)}\leftarrow\frac{(\mu^{(i)}-\bar{\mu}^{(i-1)})^{T}(\mu_{E}-\bar{\mu}^{(i-1)})}{(\mu^{(i)}-\bar{\mu}^{(i-1)})^{T}(\mu^{(i)}-\bar{\mu}^{(i-1)})}(\mu^{(i)}-\bar{\mu}^{(i-1)})$ $\triangleright$ Projection of $\mu_{E}$ onto the line through $\bar{\mu}^{(i-1)}$ and $\mu^{(i)}$
@@ II-B2 Maximum Entropy IRL
$\theta^{*}=\underset{\theta}{\operatorname{argmax}}\,\mathcal{L}(\theta):=\sum_{j=1}^{N}\log\big{(}P(\tau_{j}\mid\theta)\big{)}=\underset{\theta}{\operatorname{argmax}}\,\sum_{j=1}^{N}\log\left(\frac{\exp(r_{\tau}(\tau_{j}))}{\sum_{j=1}^{N}\exp(r_{\tau}(\tau_{j}))}\right),$ (7)
@@ III-C Implementation and Performance Guarantees for Equivocal Deception
The second term in the objective function minimizes the squared difference of the sum of occupancy measures in $\mathcal{S}_{goal}$ and $\mathcal{S}_{decoy}$. Similar to Optimization Problems 4 and 5, the weighting parameter $\beta>0$ encodes the trade-off between deception and performance. If the accumulated reward, i.e., the first term of the objective function, can be maximized while satisfying $\sum_{s\in\mathcal{S}_{goal}}\sum_{a\in\mathcal{A}}x(s,a)=\sum_{s\in\mathcal{S}_{decoy}}\sum_{a\in\mathcal{A}}x(s,a)$, then the deception constraint in Optimization Problem 6 is always satisfied. Otherwise, deception is prioritized by a larger value of $\beta$. The objective function in Optimization Problem 7 is a strongly concave quadratic function when $\beta>0$. Optimization Problem 7 has the same feasible region as Optimization Problem 3, which is known to be non-empty, and therefore a solution exists to Optimization Problem 7. Since it maximizes a strongly concave function over a convex, compact set, its solution can be computed efficiently with convex quadratic programming.
@@ IV Numerical Results
In this section, we solve Problem 2. We use our techniques for deceptive policy optimization in the setting of the moving target defense (MTD) problem from [36]. By extending the problem to a multi-agent setting, the objective is to hide a “real agent” whose task must be completed among a collection of “decoy agents” that are part of the system only to aid in the implementation of deception. We seek to do this while maintaining strong performance in the sense of still maximizing revenue as much as possible. To that end, we empirically evaluate the extent to which the policies obtained from Optimization Problems 4,  5, and 7 achieve both deception and strong performance simultaneously.  [^1] The software and data used to generate these results is available at: https://anonymous.4open.science/r/DeceptionMTD-2980..
@@ B $R-C_{B}$ $R-C_{D}-C_{B}$ $R-C_{R}$
We use three types of IRL to model the adversary: AL, MaxEnt IRL, and Deep IRL. The objective of the adversary is to identify the real agent(s) among the full collection of agent(s). After estimating the agents’ rewards using IRL, the adversary infers the real agents as the agent(s) that have higher reward in state “N” and lower reward in state “B”. If the adversary makes an incorrect estimate of the system’s rewards, then they will infer incorrect goal states, which will result in incorrect inferences of the identities of the real agent(s). Hence, to conceal the real agent from the adversary, we compute deceptive policies to influence the reward estimated by IRL.
We computed deceptive policies by solving Optimization Problems 4,  5 and 7, each with $v_{reach}=7$. We evaluated the deceptiveness of the resulting policies by comparing the weighted sum of the estimated marginal reward of each agent. For IRL’s estimate of the reward function $\tilde{r}$, we calculate the following quantity by leveraging the deception metric proposed in [18], namely
@@
for $i\in\{0,1,\dots,n+m-1\}$, where $w(y^{i})\geq 0$ are weights for each local state $y^{i}$ and where
@@
is the estimated marginal reward for agent $i$ at local state $y^{i}$. Intuitively, this metric represents the estimated importance of each agent based on marginal estimated rewards, where states with higher rewards are considered more important. A larger value of the metric $D^{i}$ for decoy agent $i$ indicates that the adversary has inferred higher rewards for that decoy agent, and thus they have been deceived.
We consider $n=1$ real agent and $m=3$ decoy agents, where Agent0 is a real agent, and Agent1, Agent2 and Agent3 are decoy agents. The resulting optimization problems are formulated over $12^{4}=20,736$ joint state-action pairs. For weights of the estimated marginal reward, we used $w(\texttt{N})=1$, $w(\texttt{T})=0.75$, $w(\texttt{E})=0.5$, and $w(\texttt{B})=0.25$ because each agent may be considered important if it has (i) higher rewards in state N and (ii) lower rewards in state B. The weighted sum of the estimated marginal reward $D^{i}$ of each agent $i$ can be interpreted as the estimated importance of agent $i$ in the system. To assess the relative importance of each agent from $D^{i}$, we convert the collection of $D^{i}$’s into a probability distribution using softmax of each $D^{i}$ with the parameter $\lambda=1000$, i.e.,
$L^{i}=\frac{\exp{(D^{i}/\lambda)}}{\sum_{i=1}^{n+m}\exp{(D^{i}/\lambda)}}.$ (22)
A larger value of $D^{i}$ will result in a larger value of $L^{i}$, and the adversary identifies the real agent as the one with the largest value of $L^{i}$.
Building on Definitions 5,  6, and 7 in the context of this problem, we say that a policy is diversionary deceptive if the adversary misidentifies the real agent in any way, it is targeted deceptive if the adversary infers that the particular choice of decoy agent is the real agent, and it is equivocal deceptive if the adversary infers that the specific decoy agent is as important as the real agent. In the next two subsections, we demonstrate the deceptiveness and performance of policies obtained from Optimization Problems 45, and 7, then validate Theorems 12, and 3 by comparing empirically obtained revenue values to theoretically predicted values.
@@ IV-A Deceptiveness
In this section, we evaluate the deceptiveness of the deceptive policy $\pi_{d}$ obtained by solving Optimization Problems 45, or 7 in adversarial settings. In the context of MTD and IRL, the adversary infers that the agent with the maximum $L^{i}$ value is the real agent. Figure 2 shows the adversary’s estimates using three different IRL algorithms: AL, MaxEnt IRL, and Deep IRL.
@@ IV-A1 Diversionary Deception
From Definition 5, a policy that is diversionary deceptive seeks to lead an adversary to infer that any non-goal state is a goal state when attempting to learn the rewards of a system. In the context of MTD, we seek to make the adversary infer that any decoy agent is the real agent. We therefore want any decoy agent to yield the highest $L^{i}$.
The first row of Figure 2 shows that this goal is attained for larger values of the deception parameter $\beta$. With $\beta\geq 0.3$, we see that Agent0, the real agent, has a lower likelihood than at least one of the decoy agents when using any of the considered IRL algorithms, implying that there exists a level of deception capable of simultaneously deceiving all of the tested algorithms. This implementation provides diversionary deception by ensuring that the adversary draws an incorrect conclusion about the identity of the real agent.
@@ IV-A2 Targeted Deception
From Definition 6, a policy that is targeted deceptive seeks to lead an adversary to a specific, user-specified false belief. In this example, we seek to lead an adversary to infer that Agent1 is the real agent, even though it is a decoy. To achieve this, we aim to manipulate the occupancy measure $\{x(s,a)\}_{s\in\mathcal{S},a\in\mathcal{A}}$ to influence the metric $L^{i}$ so that the largest value of $L^{i}$ is obtained with Agent1.
The second row of Figure 2 shows that this goal is attained for larger values of the deception parameter $\beta$. With $\beta\geq 0.2$, Agent1 has a higher likelihood than Agent0 under all three types of IRL, indicating that the adversary incorrectly infers that Agent1 is the real agent. This implementation provides targeted deception by ensuring that the adversary draws an incorrect conclusion about the identity of the real agent.

### v3

@@ Abstract
Autonomous systems are increasingly expected to operate in the presence of adversaries, though adversaries may infer sensitive information simply by observing a system. Therefore, present a deceptive sequential decision-making framework that not only conceals sensitive information, but actively misleads adversaries about it. We model autonomous systems as Markov decision processes, with adversaries using inverse reinforcement learning to recover reward functions. To counter them, we present three regularization strategies for policy synthesis problems that actively deceive an adversary about a system’s reward. “Diversionary deception” leads an adversary to draw any false conclusion about the system’s reward function. “Targeted deception” leads an adversary to draw a specific false conclusion about the system’s reward function. “Equivocal deception” leads an adversary to infer that the real reward and a false reward both explain the system’s behavior. We show how each form of deception can be implemented in policy optimization problems and analytically bound the loss in total accumulated reward induced by deception. Next, we evaluate these developments in a multi-agent setting. We show that diversionary, targeted, and equivocal deception all steer the adversary to false beliefs while still attaining a total accumulated reward that is at least $98\%$ of its optimal, non-deceptive value.
@@ I Introduction
Existing work in [22] has presented related definitions of deception and empirical results for them, but to the best of our knowledge the current paper is the first to analytically characterize these types of deception and the tradeoffs they create. In addition, this work expands the empirical results by exploring deception and performance under different algorithms of inverse reinforcement learning, ensuring their reliability in applications.
@@ II-A Preliminaries on Markov Decision Processes
$\begin{aligned} \textnormal{ subject to }\,\,&v_{\pi}(s)-\gamma\sum_{j\in\mathcal{S}}\mathcal{T}(s,a,j)v_{\pi}(j)\geq r(s,a)\textnormal{ for all }s\in\mathcal{S},a\in\mathcal{A}.\end{aligned}$
@@
$\begin{aligned} &\underset{x\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}}{\operatorname{maximize}}\quad\sum_{s\in S}\sum_{a\in A}r(s,a)x(s,a)\end{aligned}$ (3)
@@ II-B1 Apprenticeship Learning
  $\bar{\mu}^{(i)}\leftarrow\frac{(\mu^{(i)}-\bar{\mu}^{(i-1)})^{T}(\mu_{E}-\bar{\mu}^{(i-1)})}{(\mu^{(i)}-\bar{\mu}^{(i-1)})^{T}(\mu^{(i)}-\bar{\mu}^{(i-1)})}(\mu^{(i)}-\bar{\mu}^{(i-1)})$
$\triangleright$ Projection of $\mu_{E}$ onto the line through $\bar{\mu}^{(i-1)}$ and $\mu^{(i)}$
@@ II-B2 Maximum Entropy IRL
$\theta^{*}$ $=\underset{\theta}{\operatorname{argmax}}\>\mathcal{L}(\theta):=\sum_{j=1}^{N}\log\big(P(\tau_{j}\mid\theta)\big)=\underset{\theta}{\operatorname{argmax}}\sum_{j=1}^{N}\log\left(\frac{\exp(r_{\tau}(\tau_{j}))}{\sum_{j=1}^{N}\exp(r_{\tau}(\tau_{j}))}\right),$ (7)
@@ III-C Implementation and Performance Guarantees for Equivocal Deception
The second term in the objective function minimizes the squared difference of the sum of occupancy measures in $\mathcal{S}_{goal}$ and $\mathcal{S}_{decoy}$. Similar to Optimization Problems 4 and 5, the weighting parameter $\beta>0$ encodes the trade-off between deception and performance. If the accumulated reward, i.e., the first term of the objective function, can be maximized while satisfying $\sum_{s\in\mathcal{S}_{goal}}\sum_{a\in\mathcal{A}}x(s,a)=\sum_{s\in\mathcal{S}_{decoy}}\sum_{a\in\mathcal{A}}x(s,a)$, then the deception constraint in Optimization Problem 6 is always satisfied. Otherwise, deception is prioritized by a larger value of $\beta$. The objective function in Optimization Problem 7 is a strongly concave quadratic function when $\beta>0$. Optimization Problem 7 has the same feasible region as Optimization Problem 3, which is known to be non-empty, and therefore a solution exists to Optimization Problem 7. Since it maximizes a strongly concave function over a convex, compact set, its solution can be computed efficiently with convex qudratic programming.
@@ IV Numerical Results
In this section, we solve Problem 2. We use our techniques for deceptive policy optimization in the setting of the moving target defense (MTD) problem from [39], which can be generalized to real-world network defense problems, such as DDoS defense [34, 17] and protecting smart grid [14] or vehicle networks [25]. By extending the problem to a multi-agent setting, the objective is to hide “real agent(s)” whose task must be completed among a collection of “decoy agents” that are part of the system only to aid in the implementation of deception. We seek to conceal the real agent(s) while maintaining strong performance in the sense of still maximizing revenue as much as possible. To that end, we empirically evaluate the extent to which the policies obtained from Optimization Problems 4,  5, and 7 achieve both deception and strong performance simultaneously.  [^1] The software and data used to generate these results is available at: https://anonymous.4open.science/r/DeceptionMTD-2980..
@@ B $R-C_{B}$ $R-C_{D}-C_{B}$ $R-C_{R}$
We use three types of IRL to model the adversary: AL, MaxEnt IRL, and Deep IRL. The objective of the adversary is to identify the real agent(s) among the full collection of agents. After estimating the agents’ rewards using IRL, the adversary infers the real agent(s) as the agent(s) that have higher reward in state “N” and lower reward in state “B”. If the adversary makes an incorrect estimate of the system’s rewards, then they will infer incorrect goal states, which will result in incorrect inferences of the identities of the real agent(s). Hence, to conceal the real agent from the adversary, we compute deceptive policies to influence the reward estimated by IRL.
From the adversary’s perspective, they infer the real agent based on the likelihood of each agent being the real agent, computed as
$L^{i}=\frac{\exp{(D^{i}/\lambda)}}{\sum_{i=1}^{n+m}\exp{(D^{i}/\lambda)}},$ (20)
with some $D^{i}$ that measures the importance of agent $i$, and some temperature parameter $\lambda$. For IRL’s estimate of the reward function $\tilde{r}$, we calculate $D^{i}$ by leveraging the deception metric proposed in [22], namely
@@
for $i\in\{1,\dots,n+m\}$, where $w(y^{i})\geq 0$ are weights for each local state $y^{i}$ and where
@@
is the estimated marginal reward for agent $i$ at local state $y^{i}$. For weights of $D^{i}$, we used $w(\texttt{N})=1$, $w(\texttt{T})=0.75$, $w(\texttt{E})=0.5$, and $w(\texttt{B})=0.25$, reflecting that each agent may be considered important if it has (i) higher rewards in state N and (ii) lower rewards in state B. Intuitively, this metric represents the estimated importance of each agent based on the estimated rewards, where states with higher rewards are considered more important. A larger value of $D^{i}$ will result in a larger value of $L^{i}$, and the adversary identifies the real agent as the one with the largest value of $L^{i}$.
In this example, we consider 5 agents, where the number of real and decoy agents varies in each scenario. We computed deceptive policies by solving Optimization Problems 4,  5, and 7, each with $v_{reach}=7$, which are formulated over $12^{5}=248,832$ joint state-action pairs. Building on Definitions 5,  6, and 7 in the context of this problem, we say that a policy is diversionary deceptive if the adversary misidentifies the real agent in any way, it is targeted deceptive if the adversary infers that the particular choice of decoy agent is the real agent, and it is equivocal deceptive if the adversary infers that the specific decoy agent is as important as the real agent. In the next two subsections, we demonstrate the deceptiveness and performance of policies obtained from Optimization Problems 45, and 7, then validate Theorems 12, and 3 by comparing empirically obtained revenue values to theoretically predicted values.
@@ IV-A Deceptiveness
In this section, we evaluate the deceptiveness of the deceptive policy $\pi_{d}$ obtained by solving Optimization Problem 45, or 7 in adversarial settings. In the context of MTD and IRL, the adversary infers that the agent with the maximum $L^{i}$ value is the real agent. To measure the results of deception, we define the adversary’s utility under deception. After calculating $L^{i}$ based on the estimated rewards, the adversary attacks the agent that has the maximum value of $L^{i}$. As a result, the adversary gets utility $1$ if it attacked the true target, gets utility 1-1 if attacked the “target decoy agent” we selected, and gets utility 0 otherwise. In Table IIIII, and IV, we evaluate this utility of the adversary after 100 runs with different values of the parameter $\beta$. In Figure 23 and 4, we show the adversary’s estimates of $L^{i}$’s using three different IRL algorithms: AL, MaxEnt IRL, and Deep IRL.
@@ IV-A1 Diversionary Deception
From Definition 5, a policy that is diversionary deceptive seeks to lead an adversary to infer that any non-goal state is a goal state when attempting to learn the rewards of a system. In the context of MTD, we seek to make the adversary infer that any decoy agent is the real agent. We therefore want any decoy agent to yield the highest $L^{i}$. To verify this goal, we considered 2 scenarios that have different ratios of real and decoy agents, where we have one real and four agents decoy in scenario 1, and two real and three decoy agents in scenario 2.
Figure 2: The adversary’s estimate of $L^{i}$ of each agent under diversionary deception with three types of IRL. Each plot shows results averaged over $100$ runs. We used $\textit{number of trajectories}=500$ and $\textit{length of trajectories}=500$ for all simulations, $\textit{epoch}=20$, $\textit{learning rate}=0.5$ for MaxEnt IRL, and $\textit{epoch}=50$, $\textit{learning rate}=0.1$ and two fully connected hidden layers with dimension $64$ and $32$ for Deep IRL. The blue line(s) represents the real agent(s), while the gray lines represent decoy agents. When $\beta=0$, no deception is implemented, and the adversary correctly infers that the real agent(s) is the most important. However, with $\beta\geq 0.4$, the adversary’s inference of the most important agent will be one of the decoy agents in average, confirming that the policy satisfies the definition of diversionary deception.
Deception Parameter $\beta$ 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
Scenario 1 Real 100 100 100 100 0 0 5 2 0 42 22
Decoy 1 0 0 0 0 34 17 26 18 18 30 28
Decoy 2 0 0 0 0 21 13 22 17 22 5 12
Decoy 3 0 0 0 0 20 8 23 30 31 3 8
Decoy 4 0 0 0 0 25 62 24 33 29 20 30
Adversary’s Utility 100 100 100 100 0 0 5 2 0 42 22
Scenario 2 Real 1 100 45 19 16 0 0 16 7 3 1 1
Real 2 0 47 14 18 0 0 12 7 2 3 3
Decoy 1 0 2 23 25 35 36 31 41 44 33 32
Decoy 2 0 3 21 20 38 28 20 26 35 34 34
Decoy 3 0 3 23 21 27 36 21 19 16 29 30
Adversary’s Utility 100 92 33 34 0 0 28 14 5 4 4
TABLE II: Adversary’s estimation of the real agent and total utility when using Deep IRL under diversionary deception
Figure 2 shows that this goal is attained in both scenarios for larger values of the deception parameter $\beta$. With $\beta\geq 0.4$, we see that the real agent has a lower likelihood than at least one of the decoy agents when using any of the considered IRL algorithms, implying that there exists a level of deception capable of simultaneously deceiving all of the tested algorithms. Table II shows the estimation of the real agent and the resulting utility when its using Deep IRL. In both scenarios, with $\beta=0$ (when there is no deception), the adversary always find out the real agent(s) and successfully attacks them. However, with $\beta\simeq 0.4$, the adversary never attacks the real agent(s), while larger values of $\beta$ tends to show mixed likelihood, making all agents indistinguishable in their importance. This implementation provides diversionary deception with $\beta\geq 0.4$ by ensuring that the adversary draws an incorrect conclusion about the identity of the real agent(s).
@@ IV-A2 Targeted Deception
From Definition 6, a policy that is targeted deceptive seeks to lead an adversary to a specific, user-specified false belief. In the context of MTD, we seek to lead an adversary to infer that one of the target decoy agents is the real agent. To verify this goal, we consider 2 scenarios that have different ratios of real and target decoy agent(s), where we have one real and one target decoy agents in scenario 1, and one real and two target decoy agents in scenario 2. To achieve targeted deception, we aim to manipulate the occupancy measure $\{x(s,a)\}_{s\in\mathcal{S},a\in\mathcal{A}}$ to influence the metric $L^{i}$ so that the largest value of $L^{i}$ is obtained with the target decoy agent(s).
Figure 3: The adversary’s estimate of $L^{i}$ of each agent under targeted deception with three types of IRL. Each plot shows results averaged over $100$ runs. We used $\textit{number of trajectories}=500$ and $\textit{length of trajectories}=500$ for all simulations, $\textit{epoch}=20$, $\textit{learning rate}=0.5$ for MaxEnt IRL, and $\textit{epoch}=50$, $\textit{learning rate}=0.1$ and two fully connected hidden layers with dimension $64$ and $32$ for Deep IRL. The blue line represents the real agent, the red line(s) represents the target decoy agent(s), and the gray lines represent decoy agents. When $\beta=0$, no deception is implemented, and the adversary correctly infers that the real agent is the most important. With $\beta\geq 0.02$ the likelihood of the target decoy agent(s) is greater than that of the real agent, implying that deception misleads the adversary into believing that the target decoy agent(s) is the real one. Thus, the policy satisfies the definition of targeted deception.
Deception Parameter $\beta$ 0 0.02 0.04 0.06 0.08 0.1 0.12 0.14 0.16 0.18 0.2
Scenario 1 Real 99 0 0 0 0 0 0 0 0 0 0
Target decoy 0 100 100 100 100 100 100 100 100 100 100
Decoy 1 0 0 0 0 0 0 0 0 0 0 0
Decoy 2 1 0 0 0 0 0 0 0 0 0 0
Decoy 3 0 0 0 0 0 0 0 0 0 0 0
Adversary’s Utility 99 -100 -100 -100 -100 -100 -100 -100 -100 -100 -100
Scenario 2 Real 100 0 0 0 0 0 0 0 0 0 0
Target decoy 1 0 47 50 52 54 54 58 55 53 46 56
Target decoy 2 0 53 50 48 46 46 42 45 47 54 44
Decoy 1 0 0 0 0 0 0 0 0 0 0 0
Decoy 2 0 0 0 0 0 0 0 0 0 0 0
Adversary’s Utility 100 -100 -100 -100 -100 -100 -100 -100 -100 -100 -100
TABLE III: Adversary’s estimation of the real agent and total utility when using Deep IRL under targeted deception
Figure 3 shows that targeted deception is attained for larger values of $\beta$. With $\beta\geq 0.02$, the target decoy agent(s) has a larger $L^{i}$ than the real agent, under all three types of IRL. This indicates that the adversary incorrectly infers that the target decoy agent(s) is the real agent. Table III shows that the adversary’s estimation of the real agent is always one of the target decoy agents with $\beta\geq 0.02$, resulting in a substantial decrease in the adversary’s utility. This implementation provides targeted deception by ensuring that the adversary draws a specific incorrect conclusion about the identity of the real agent.